# DATABench: Evaluating Dataset Auditing in Deep Learning from an Adversarial Perspective

### v1

@@ Abstract
The widespread application of Deep Learning across diverse domains hinges critically on the quality and composition of training datasets. However, the common lack of disclosure regarding their usage raises significant privacy and copyright concerns. Dataset auditing techniques, which aim to determine if a specific dataset was used to train a given suspicious model, provide promising solutions to addressing these transparency gaps. While prior work has developed various auditing methods, their resilience against dedicated adversarial attacks remains largely unexplored. To bridge the gap, this paper initiates a comprehensive study evaluating dataset auditing from an adversarial perspective. We start with introducing a novel taxonomy, classifying existing methods based on their reliance on internal features (IF) (inherent to the data) versus external features (EF) (artificially introduced for auditing). Subsequently, we formulate two primary attack types: evasion attacks, designed to conceal the use of a dataset, and forgery attacks, intending to falsely implicate an unused dataset. Building on the understanding of existing methods and attack objectives, we further propose systematic attack strategies: decoupling, removal, and detection for evasion; adversarial example-based methods for forgery. These formulations and strategies lead to our new benchmark, DATABench, comprising 17 evasion attacks, 5 forgery attacks, and 9 representative auditing methods. Extensive evaluations using DATABench reveal that none of the evaluated auditing methods are sufficiently robust or distinctive under adversarial settings. These findings underscore the urgent need for developing a more secure and reliable dataset auditing method capable of withstanding sophisticated adversarial manipulation. Code is available at https://github.com/shaoshuo-ss/DATABench.
I Introduction
The rapid advancement of Deep Learning (DL) has propelled it into widespread applications across various domains and real-world systems, such as computer vision [32], artificial intelligence (AI) chatbots [28], and self-driving vehicles [10]. Currently, DL models primarily make predictions based on their knowledge learned from a dataset. As such, its success highly relies on and is inseparable from high-quality datasets [67, 45, 50]. As a critical resource and fundamental component of DL, datasets largely determine the performance of the models trained on them. However, developers of DL models and applications usually do not disclose the training datasets they use. This lack of transparency gives rise to potential privacy and copyright concerns. For instance, training data may include unauthorized private information. Commercially deployed models may have been trained using open-source copyrighted datasets that prohibit commercial use. Additionally, many existing data protection regulations, such as GDPR, have granted data owners the right to be informed of their data usage. Consequently, auditing the datasets used in DL models has become a critical issue for safeguarding data privacy and copyright [41, 18].
@@ I Introduction
Figure 1: The general pipeline of dataset auditing. In the preparation stage, the data owner collects and publishes a dataset. In the exploitation stage, an adversary may copy or steal the dataset and train a model for unauthorized purposes. Finally, in the auditing stage, the data owner audits the suspicious model and verifies whether it is trained on the protected dataset.
@@
Dataset auditing [^1] Existing literature may use different terminologies, such as dataset ownership verification [49], dataset copyright auditing [18], or data-use auditing [41]. In this paper, we standardize the term as dataset auditing in DL. in DL [49, 18, 42] refers to determining whether a specific dataset is used to train a suspicious model. Typically, as shown in Figure 1, there are three stages involved in dataset auditing [53]. In the first stage, the data owner collects some data from various resources. Subsequently, he processes the dataset ($e.g.$, data annotation) to make it suitable for usage or release. Then, in the second stage, an adversary may get access to the dataset through copying or stealing and utilize it to train a DL model without authorization. Third, if a third-party DL model (dubbed a ‘suspicious model’) is suspected of using a protected dataset unauthorizedly, dataset auditing techniques will be adopted to verify it. However, there exist several challenges to achieving a reliable and accurate dataset auditing. (1) It is hard to uncover the traces of a dataset only from the parameters, intermediate results, or even black-box access to a trained DL model. These revealed clues may not be directly associated with the data. (2) The adversary can use various different model architectures and training techniques to produce the model. The diversity and the training-agnostic situation make it difficult to achieve robust and distinctive dataset auditing. (3) The adversary may intentionally process the data or model to compromise the auditing procedure, challenging the robustness of data auditing solutions. As such, achieving dataset auditing is a non-trivial issue in practice [18, 71, 13].
@@
Although some existing works have proposed promising solutions to dataset auditing [79, 29, 60], there is still a significant gap between how the robustness and distinctiveness of existing works are evaluated (often in controlled environments) and how they perform in real-world practice. Existing dataset auditing research primarily considers naive operations and techniques in DL under constrained settings, such as training with models of different architectures, fine-tuning, and pruning, as potential adversarial vectors [49, 18, 31]. These works lack the consideration of a ‘clever’ adversary who may design specified and adaptive attacks against dataset auditing. It is challenging to conduct an exhaustive evaluation due to the complexity of training a model in DL and the diversity of attack surfaces. Consequently, the following critical question remains unanswered: are existing dataset auditing methods reliable under adversarial settings?
@@
We develop the Dataset AudiTing Attack Benchmark, DATABench, based on the proposed attack formulations and strategies. Specifically, we implement 17 evasion attacks and 5 forgery attacks, and evaluate 9 representative dataset auditing methods on the proposed benchmark. The empirical results reveal that none of the evaluated auditing methods exhibit sufficient robustness or distinctiveness under adversarial settings. We surprisingly find that even simple attacks can significantly compromise the effectiveness of existing methods. The results highlight an urgent need for the development of a more reliable dataset auditing method. Based on our findings, we also point out several potential future directions that can provide guidance to develop more reliable dataset auditing methods.
@@
DATABench also provides an extensible open-source toolbox to facilitate future research and development in this crucial area. This toolbox is designed to streamline the process for researchers to easily develop new dataset auditing techniques, implement additional attacks, and systematically evaluate their proposed methods against the comprehensive and standardized attack suite provided within our unified benchmark.
II Background
II-A Deep Learning
@@
Deep Learning (DL) has become the most famous and widely-used technique to develop an artificial intelligence (AI) application during the last decade [32, 67, 33], especially with the growth of large foundation models [28]. DL primarily leverages deep neural networks (DNNs) as the core model architecture to achieve a wide range of functions.
@@
During the development of a DL model $\mathcal{M}$, the model developer needs to first collect a dataset $\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{|\mathcal{D}|}$. The dataset $\mathcal{D}$ is related to the developer-specified DL task. The dataset consists of various input samples $\bm{x}_{i}$ and their corresponding label $\bm{y}_{i}$. Subsequently, the developer designs a loss function $\mathcal{L}(\cdot)$ and optimizes the model $\mathcal{M}$ according to the following optimization problem:
^$\widehat{\mathcal{M}}=\arg\min_{\mathcal{M}}\sum_{i=1}^{k}\mathcal{L}(\mathcal{M}(\bm{x}_{i}),\bm{y}_{i}).$ (1)
@@
Currently, most of the DL techniques and models are data-centric and their developments highly rely on a high-quality dataset [67]. This reliance, however, raises significant concerns regarding data privacy and copyright. For instance, many DL models are trained on large-scale datasets that may contain sensitive or proprietary information, leading to potential violations of user privacy and intellectual property rights [6, 18]. Consequently, data protection and building high-quality datasets are equally important in practice.
II-B Data Protection
Traditional Data Watermarking. Traditional data watermarking techniques help track ownership and usage of data by embedding hidden markers ($i.e.$, watermark) into the data [43, 87]. The data owner can extract the watermark from the data and prove their ownership. Multiple solutions to embedding a watermark into data have been proposed, such as embedding watermarks into the frequency domain or utilizing DL models [8, 4]. Recently, some works have also explored using data watermarking to mark AI-generated content and models [86, 44, 69]. However, traditional data watermarking techniques may not be feasible for dataset auditing since the watermark in the data can hardly be inherited by the DL models and their owners will not disclose their training details (especially their training data).
III Dataset Auditing in Deep Learning
In this section, we present the formal definition, our novel taxonomy, and the objectives of dataset auditing methods. In this paper, we focus on dataset auditing for image classification models, since most existing methods work on them, and they usually cannot be directly adapted to other modalities and tasks. Further discussion about dataset auditing for generative models can be found in Section VI-B.
III-A Towards a Formal Definition of Dataset Auditing
@@
Before defining dataset auditing, we hereby first introduce the general threat model of dataset auditing in DL [49, 18]. Specifically, there are two parties involved in the threat model, the data owner and the adversary [^2] Some existing literature [56, 76, 18] also introduces a trustworthy third-party auditor (or verifier) in their threat model to conduct the audit. However, since the auditor does not affect the technical design of dataset auditing methods, we exclude it to avoid potential confusion.. The data owner wants to collect, store, and publish a DL dataset. However, there exists the risk that an adversary may unlawfully copy or steal the dataset and train his own model for unauthorized usage. Such a misbehavior may compromise the privacy or intellectual property rights of the data owner. As such, the data owner seeks an effective dataset auditing method to confirm whether the suspicious model of the adversary is trained on his dataset. Now, we formally define dataset auditing as follows.
@@
Notably, if the number of samples in $\mathcal{D}$ is $1$ ($i.e.$, $|\mathcal{D}|=1$), the definition of dataset auditing is similar to membership inference [6, 65, 74]. As such, existing membership inference techniques can be easily adapted to achieve dataset auditing, which will be further discussed in Section III-B.
@@
    Preparation Stage. The data owner collects the data and organizes it into a raw dataset $\mathcal{D}$. Then, before publishing the dataset, the data owner can also apply a processing function $\mathcal{P}(\cdot)$: $\widehat{\mathcal{D}}=\mathcal{P}(\mathcal{D},\mathcal{I}),$ where $\mathcal{I}$ is the auxiliary information needed in the processing function. $\mathcal{P}(\cdot)$ may be used to add specific perturbations or watermarks to the dataset, or just keep $\widehat{\mathcal{D}}=\mathcal{D}$. Then, the data owner publishes the dataset $\widehat{\mathcal{D}}$.
@@
    Exploitation Stage. After the data owner publishes the dataset $\widehat{\mathcal{D}}$, an adversary may steal or copy $\widehat{\mathcal{D}}$ to train his own model $\mathcal{M}$ for unauthorized purposes. He can adopt any training techniques and conduct certain operations to intentionally compromise the dataset auditing mechanism.
@@
    Auditing Stage. If the data owner finds a model [^3] Following prior works [18, 49, 41, 53], we do not focus on how to identify these ‘suspicious models’, which may be discovered through non-technical ways. For instance, if a model’s task closely aligns with that of a dataset, the model may be regarded as a ‘suspicious model’. that is suspected of unauthorizedly using the dataset $\widehat{\mathcal{D}}$, he adopts a dataset auditing method $\mathcal{A}$ to judge whether the misbehavior exists. If $\mathcal{A}(\widehat{\mathcal{D}},\mathcal{M},\mathcal{I})$ outputs $1$, the adversary can be accused of unauthorizedly using the dataset and infringing the copyright or privacy.
@@
  • In the preparation stage, the data owner has full control of this stage and the dataset. He can determine what preprocessing steps $\mathcal{P}(\cdot)$ to apply to the dataset and which portions of the data to release.
  • In the exploitation stage, the data owner has no knowledge of the operations of the adversary. He also cannot interfere with the training of the adversary’s model.
@@
In the preparation stage, if the data owner does not need to alter the raw dataset, $i.e.$, $\mathcal{D}=\widehat{\mathcal{D}}$, the dataset auditing method is a non-intrusive method. Otherwise, it is an intrusive method.
@@
In this paper, we pay more attention to black-box dataset auditing since it is more practical in real-world scenarios.
III-B Dataset Auditing Taxonomy
@@ III-C Design Objectives
  • Robustness: The dataset auditing method should not be significantly affected by an adversary’s attacks $i.e.$, still outputs $1$ when the model $\mathcal{M}$ is indeed trained on $\widehat{\mathcal{D}}$.
  • Distinctiveness: The dataset auditing method should not identify other unused datasets as being used for training, $i.e.$, still outputs 00 when $\mathcal{M}$ is not trained on $\mathcal{D}_{u}$.
@@
IV Attack Methodology
As introduced in Section III-C, robustness and distinctiveness are two fundamental and important properties of a practical dataset auditing method. However, there is still a significant gap between the constrained settings of existing works and practice for evaluating these two properties. Existing dataset auditing research [49, 18, 31] primarily considers common yet simple operations in DL and lacks dedicated attack evaluation. To bridge this gap, we propose a unified and comprehensive formulation and methodology of attacks against dataset auditing. We also present systematic attack strategies and implementations of these attacks. All these elements lead to the development of our Dataset Auditing Attack Benchmark, DATABench.
IV-A Attack Formulation
@@
where $\mathcal{I}$ is the auxiliary information required in the process of dataset auditing. The output 00 indicates that $\widetilde{\mathcal{M}}$ is not trained on $\mathcal{D}$. Meanwhile, $\widetilde{\mathcal{M}}$ needs to have a similar utility with the model directly trained on $\mathcal{D}$.
@@ IV-B Evasion Attacks
Attack Strategy. According to Section III-B, existing dataset auditing methods can be classified into internal feature (IF) based methods and external feature (EF) based methods. This taxonomy can facilitate the design of attacks, as follows.
@@
  • Decoupling: To reduce the association between the trained model and the dataset, the adversary can adopt operations to make the samples in the dataset less similar to the original ones. As such, we design and implement data augmentation and data synthesis as two approaches to achieve such a goal. Data augmentation applies transformations such as cropping and flipping to increase the diversity of training data [78]. Data synthesis is an advanced technique that utilizes the capability of generative models. It first trains a generative model ($e.g.$, a diffusion model [36]) on the dataset and then uses the model to generate a synthetic dataset. The model is trained on the synthetic dataset instead of the original one.
@@
  • Removal: To eliminate potential abnormal behaviors during training, we consider adversarial training and training-phase backdoor defense as two representative removal attacks. Adversarial training [80] improves the prediction robustness of the model by including adversarial examples (AE) during the training process. AEs are inputs intentionally perturbed to lead to misclassification; Training-phase backdoor defense [26, 40] aims to mitigate hidden external features in the training data by trying not to learn from them. Both ways may remove the designated behaviors and evade dataset auditing.
@@
  • Decoupling: In this phase, the adversary can also apply DP mechanisms to add noise during the model inference [19]. As such, we implement noisy output and noisy feature that perturb the outputs and intermediate results of the model, respectively, during inference to evade dataset auditing.
  • Removal: Since the model parameters are fixed, the adversary cannot remove the external features from the model. However, it can still hide the external behaviors by changing inference mechanisms. Randomized smoothing (RS) and transformation-based backdoor defense are two typical methods. Specifically, RS [15] adds noise to the inputs and calculates the aggregated results of the outputs on noisy inputs; Transformation-based backdoor defense [52, 12] ($e.g.$, reprogramming) neutralizes the designated behaviors by applying input transformations.
  • Detection: This type of attack aims to detect the trigger samples used for dataset auditing. Since the trigger samples tend to be different from normal samples, we implement traditional outlier detection based on the features of inputs and advanced backdoor sample detection [37, 30]. If these methods identify the input samples as an outlier or backdoor sample, the adversary can refuse to reply to evade dataset auditing.
IV-C Forgery Attacks
Attack Strategy. The goal of the forgery attack is to craft some samples to falsely claim that an independent model $\mathcal{M}_{u}$ is trained on these samples. Assuming that the adversary has an auxiliary dataset $\mathcal{D}_{a}$, he can optimize a set of perturbations =$\mathcal{R}=\{\bm{r}_{i}\}_{i}^{|\mathcal{D}_{a}|}$ to make the dataset auditing method output $1$. Ideally, this can be formalized as an optimization problem:

### v2

@@ Abstract
The widespread application of Deep Learning across diverse domains hinges critically on the quality and composition of training datasets. However, the common lack of disclosure regarding their usage raises significant privacy and copyright concerns. Dataset auditing, which aims to determine if a specific dataset was used to train a given suspicious model, provides promising solutions to addressing these transparency gaps. While prior work has developed various auditing methods, their resilience against dedicated adversarial attacks remains largely unexplored. To bridge the gap, this paper initiates a comprehensive study evaluating dataset auditing from an adversarial perspective. We start with introducing a novel taxonomy, classifying existing methods based on their reliance on internal features (IF) (inherent to the data) versus external features (EF) (artificially introduced for auditing). Subsequently, we formulate two attack types: evasion attacks, designed to conceal the use of a dataset, and forgery attacks, intending to falsely implicate an unused dataset. Building on the understanding of existing methods and attack objectives, we further propose systematic attack strategies: decoupling, removal, and detection for evasion; adversarial example-based methods for forgery. These formulations and strategies lead to our new benchmark, DATABench, comprising 17 evasion attacks, 5 forgery attacks, and 9 representative auditing methods. Extensive evaluations using DATABench reveal that none of the evaluated auditing methods are sufficiently robust or distinctive under adversarial settings. These findings underscore the urgent need for developing a more secure and reliable dataset auditing method capable of withstanding sophisticated adversarial manipulation.
1 Introduction
@@
Figure 1: The pipeline of dataset auditing. In the preparation stage, the data owner collects and publishes a dataset. In the exploitation stage, an adversary may copy or steal the dataset and train a model for unauthorized purposes. Finally, in the auditing stage, the data owner audits the suspicious model and verifies whether it is trained on the protected dataset.
The rapid advancement of Deep Learning (DL) has propelled it into widespread applications across various real-world systems, such as computer vision [34], artificial intelligence (AI) chatbots [30], and self-driving vehicles [11]. Currently, DL models primarily make predictions based on their knowledge learned from a dataset. As such, its success highly relies on and is inseparable from high-quality datasets [68, 46]. As a critical resource and fundamental component of DL, datasets largely determine the performance of the models trained on them. However, developers of DL models and applications usually do not disclose the training datasets they use. This lack of transparency gives rise to potential privacy and copyright concerns. For instance, training data may include unauthorized private information. Commercially deployed models may have been trained using copyrighted datasets that prohibit commercial use. Additionally, many existing data protection regulations, such as GDPR, have granted data owners the right to be informed of their data usage. Consequently, auditing the datasets used in DL models has become a critical issue for safeguarding data privacy and copyright [42, 20].
@@
Dataset auditing [^1] Existing literature may use different terminologies, such as dataset ownership verification [51], dataset copyright auditing [20], or data-use auditing [42]. In this paper, we standardize the term as dataset auditing in DL. in DL [51, 20, 43] refers to determining whether a specific dataset is used to train a suspicious model. Typically, as shown in Figure 1, there are three stages involved in dataset auditing [54]. In the first stage, the data owner collects some data from various resources. Subsequently, they process the dataset ($e.g.$, data annotation) to make it suitable for usage or release. Then, in the second stage, an adversary may get access to the dataset through copying or stealing and utilize it to train a DL model without authorization. Third, if a third-party DL model (dubbed a ‘suspicious model’) is suspected of using a protected dataset unauthorizedly, dataset auditing techniques will be adopted to verify it. However, there exist several challenges to achieving a reliable and accurate dataset auditing. (1) It is hard to uncover the traces of a dataset only from the parameters, intermediate results, or even black-box access to a trained DL model. These revealed clues may not be directly associated with the data. (2) The adversary can use various different model architectures and training techniques to produce the model. The diversity and the training-agnostic situation make it difficult to achieve robust and distinctive dataset auditing. (3) The adversary may intentionally process the data or model to compromise the auditing procedure, challenging the robustness of data auditing solutions. As such, achieving dataset auditing is a non-trivial issue in practice [20, 71, 14].
@@
Although some existing works have proposed promising solutions to dataset auditing [79, 31, 61], there is still a significant gap between how the robustness and distinctiveness of existing works are evaluated (often in controlled environments) and how they perform in real-world practice. Existing dataset auditing research primarily considers naive operations and techniques in DL under constrained settings, such as training with models of different architectures, fine-tuning, and pruning, as potential adversarial vectors [51, 20, 33]. These works lack the consideration of a ‘clever’ adversary who may design specified and adaptive attacks against dataset auditing. It is challenging to conduct an exhaustive evaluation due to the complexity of training a model in DL and the diversity of attack surfaces. Consequently, the following critical question remains unanswered: are existing dataset auditing methods reliable under adversarial settings in practice?
@@
We develop the Dataset AudiTing Attack Benchmark, DATABench, based on the proposed attack formulations and strategies. Specifically, we implement 17 evasion attacks and 5 forgery attacks, and evaluate 9 representative dataset auditing methods on the proposed benchmark. The empirical results reveal that none of the evaluated auditing methods exhibit sufficient robustness or distinctiveness under adversarial settings. We surprisingly find that even simple attacks can significantly compromise the effectiveness of existing methods. The results highlight an urgent need for the development of a more reliable dataset auditing method.
@@
DATABench also provides an extensible open-source toolbox to facilitate future research and development in this crucial area. This toolbox is designed to streamline the process to easily develop new dataset auditing techniques, implement additional attacks, and systematically evaluate their proposed methods against the comprehensive and standardized attack suite provided within our unified benchmark.
2 Dataset Auditing in Deep Learning
2.1 Preliminaries
@@
Deep Learning. Deep Learning (DL) has become the most famous and widely-used technique to develop an artificial intelligence (AI) application during the last decade [34, 68], especially with the growth of large foundation models [30]. DL primarily leverages deep neural networks (DNNs) as the core model architecture to achieve a wide range of functions.
@@
While developing a DL model $\mathcal{M}$, the model developer needs to first collect a dataset $\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{|\mathcal{D}|}$. $\mathcal{D}$ is related to the developer-specified DL task and consists of various input samples $\bm{x}_{i}$ and their corresponding label $\bm{y}_{i}$. Subsequently, the developer designs a loss function $\mathcal{L}(\cdot)$ and optimizes the model $\mathcal{M}$ according to the following optimization problem:
$\widehat{\mathcal{M}}=\arg\min_{\mathcal{M}}\sum_{i=1}^{k}\mathcal{L}(\mathcal{M}(\bm{x}_{i}),\bm{y}_{i}).$ (1)
@@
Privacy and Copyright Concern about DL Datasets. Currently, most of the DL techniques and models are data-centric, and their developments highly rely on a high-quality dataset [68]. This reliance, however, raises significant concerns regarding data privacy and copyright. For instance, many DL models are trained on large-scale datasets that may contain sensitive or proprietary information, leading to potential violations of user privacy and intellectual property rights [7, 20]. However, classical data protection techniques, such as encryption [65], differential privacy [1], traditional data watermarking [87], and unlearnable data [40], are not sufficient for resolving this issue (details in Appendix J). Consequently, dataset auditing, as a technique that has negligible influence on the data utility and can protect the privacy and copyright in DL scenarios, has become an emerging yet important solution to data protection [48, 20].
2.2 Towards a Formal Definition of Dataset Auditing in DL
@@
Before defining dataset auditing, we hereby first introduce the general threat model of dataset auditing in DL [51, 20]. Specifically, there are two parties involved in the threat model, the data owner and the adversary [^2] Some existing literature [57, 76, 20] also introduces a trustworthy third-party auditor (or verifier) in their threat model to conduct the audit. However, since the auditor does not affect the technical design of dataset auditing methods, we exclude it to avoid potential confusion.. The data owner wants to collect, store, and publish a DL dataset. However, there exists the risk that an adversary may unlawfully copy or steal the dataset and train their own model for unauthorized usage. Such a misbehavior may compromise the privacy or intellectual property rights of the data owner. As such, the data owner seeks an effective dataset auditing method to confirm whether the suspicious model of the adversary is trained on their dataset. We formally define dataset auditing as follows.
@@
Notably, if the number of samples in $\mathcal{D}$ is 11 ($i.e.$, $|\mathcal{D}|=1$), the definition of dataset auditing is similar to membership inference [7, 66, 74]. As such, existing membership inference techniques can be easily adapted to achieve dataset auditing, which will be further discussed in Section 2.3.
@@
    Preparation Stage. The data owner collects the data and organizes it into a raw dataset $\mathcal{D}$. Then, before publishing the dataset, the data owner can also apply a processing function $\mathcal{P}(\cdot)$, as follows.
    $\widehat{\mathcal{D}}=\mathcal{P}(\mathcal{D},\mathcal{I}),$ (3)
    where $\mathcal{I}$ is the auxiliary information needed in the processing function. $\mathcal{P}(\cdot)$ may be used to add specific perturbations or watermarks to the dataset, or just keep $\widehat{\mathcal{D}}=\mathcal{D}$. Then, the data owner publishes the dataset $\widehat{\mathcal{D}}$.
@@
    Exploitation Stage. After the data owner publishes the dataset $\widehat{\mathcal{D}}$, an adversary may steal or copy $\widehat{\mathcal{D}}$ to train their own model $\mathcal{M}$ for unauthorized purposes. They can adopt any training techniques and conduct certain operations to intentionally compromise the dataset auditing mechanism.
@@
[…]
@@
  • In the preparation stage, the data owner has full control of this stage and the dataset. They can determine what preprocessing steps $\mathcal{P}(\cdot)$ to apply to the dataset and which portions of the data to release.
  • In the exploitation stage, the data owner has no knowledge of the operations of the adversary. They also cannot interfere with the training of the adversary’s model.
@@
In the preparation stage, if the data owner does not need to alter the raw dataset, $i.e.$, $\mathcal{D}=\widehat{\mathcal{D}}$, the dataset auditing method is a non-intrusive method. Otherwise, it is an intrusive one.
@@
2.3 Dataset Auditing Taxonomy
@@ 2.4 Design Objectives
  • Robustness: The dataset auditing method should not be significantly affected by an adversary’s attacks.
  • Distinctiveness: The dataset auditing method should not identify other unused datasets as being used for training.
@@
3 Attack Methodology
As introduced in Section 2.4, robustness and distinctiveness are two fundamental properties of a practical dataset auditing method. However, there is still a significant gap between the constrained settings of existing works and practice for evaluating these two properties. Existing research [51, 20, 33] primarily considers common yet simple operations in DL and lacks dedicated attack evaluation. To bridge this gap, we propose a unified and comprehensive formulation and methodology of attacks against dataset auditing. We also present systematic attack strategies and implementations of these attacks.
3.1 Attack Formulation
@@
where $\mathcal{I}$ is the auxiliary information required in the process of dataset auditing. The output 0 indicates that $\widetilde{\mathcal{M}}$ is not trained on $\mathcal{D}$. Meanwhile, $\widetilde{\mathcal{M}}$ needs to have a similar utility with the model directly trained on $\mathcal{D}$.
@@ 3.2 Evasion Attacks
Attack Strategy. According to Section 2.3, existing dataset auditing methods can be classified into internal feature (IF) based methods and external feature (EF) based methods. This taxonomy can facilitate the design of attacks, as follows.
@@
  • Decoupling: To reduce the association between the trained model and the dataset, the adversary can adopt operations to make the samples in the dataset less similar to the original ones. As such, data augmentation and data synthesis are two approaches to achieve such a goal. Data augmentation applies transformations such as cropping and flipping to increase the diversity of training data [78]. Data synthesis is an advanced technique that utilizes the capability of generative models [49, 18]. It first trains a generative model ($e.g.$, a diffusion model [37]) on the dataset and then uses the model to generate a synthetic dataset. The model is trained on the synthetic dataset instead of the original one.
@@
  • Removal: To eliminate potential abnormal behaviors during training, we consider adversarial training and training-phase backdoor defense as two representative removal attacks. Adversarial training [80] improves the prediction robustness of the model by including adversarial examples (AE) during the training process. AEs are inputs intentionally perturbed to lead to misclassification. Training-phase backdoor defense [28, 41] aims to mitigate hidden external features in the training data by trying not to learn from them. Both ways may remove the designated behaviors and evade dataset auditing.
@@
  • Decoupling: In this phase, the adversary can also add noise during the model inference. As such, we implement noisy output [20] and noisy feature [21] that perturb the outputs and intermediate results of the model, respectively, during inference to evade dataset auditing.
  • Removal: Since the model parameters are fixed, the adversary cannot remove the external features from the model. However, it can still hide the external behaviors by changing inference mechanisms. Randomized smoothing (RS) and transformation-based backdoor defense are two typical methods. Specifically, RS [16] adds noise to the inputs and calculates the aggregated results of the outputs on noisy inputs. Transformation-based backdoor defense [53] ($e.g.$, reprogramming [13]) neutralizes the designated behaviors by applying input transformations.
  • Detection: This type of attack aims to detect the trigger samples used for dataset auditing. Since the trigger samples tend to be different from normal samples, we implement traditional outlier detection [6] based on the features of inputs and advanced backdoor sample detection [38, 32]. If these methods identify the input samples as an outlier or backdoor sample, the adversary can refuse to reply to evade dataset auditing.
3.3 Forgery Attacks
Attack Strategy. The goal of the forgery attack is to craft samples to falsely claim that an independent model $\mathcal{M}_{u}$ is trained on these samples. Assuming that the adversary has an auxiliary dataset $\mathcal{D}_{a}$, they can optimize a set of perturbations $\mathcal{R}=\{\bm{r}_{i}\}_{i}^{|\mathcal{D}_{a}|}$ to make the dataset auditing method output $1$. Ideally, this can be formalized as an optimization problem: